OpenStudy (anonymous):

i will medal and fan anyone who can help me What is the x-coordinate of the solution to the system of equations? 2x + 6y = 8 4x – 5y = 33

OpenStudy (anonymous):

what you would do first is figure out what x is and then you will plug that into both equations

OpenStudy (anonymous):

so i solve one equation to figure out what x is and then plug it into the other equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok let me try the first one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for the first one i got x=−3y+4 i am really confused @kgod

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

take your time

OpenStudy (anonymous):

im gonna solve it and send it to you

OpenStudy (anonymous):

now plug in -3y+4 for x in both of the equations

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

@kgod was i suppose to get a fraction

OpenStudy (anonymous):

you can get a fraction that is fine

OpenStudy (anonymous):

ok i got 2/3 for the first one

OpenStudy (anonymous):

now plug it into the other equation

OpenStudy (anonymous):

x=7 and y=-1

OpenStudy (anonymous):

thanks Mrs.Allison how did you get that ?

OpenStudy (anonymous):

@Mrs.Allison is correct

OpenStudy (anonymous):

ill send you in a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I became a fan of you both k

OpenStudy (anonymous):

same.. thanks very much

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

in the first equation you isolate "2x", so it goes like this: 2x=8-6y now you place that in the place of x in the other equation, note that the other equation is "4x" so you have to multiply this expression by 2

OpenStudy (anonymous):

@MiggityMiggityMac I gave you the best response

OpenStudy (anonymous):

then it goes like: 2(8-6y)-5y=33

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

solving this you get: 16-12y-15y=33 -17y=33-16 -17y=17 y=-1

OpenStudy (anonymous):

and ok @Mrs.Allison

OpenStudy (anonymous):

i think i need to review this stuff some more

OpenStudy (anonymous):

why do you isolate 2x?

OpenStudy (anonymous):

you need thes rest or you got it?

OpenStudy (anonymous):

@Mrs.Allison woudnt you have to divide2x=8-6 to get x not 2x

OpenStudy (anonymous):

do you do that for the second one too?

OpenStudy (anonymous):

afk

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

i could isolate y, its your choice, you just have to isolate one of the letters. Also, you coould isolate just x then it would be: x= 8-6y/2

OpenStudy (anonymous):

(all divided by 2)

OpenStudy (anonymous):

and that would come out to be x = -3y +4

OpenStudy (anonymous):

so, after you found out y=-1, you go back to that first equation and place the -1 in place of y and youll find x

OpenStudy (anonymous):

yes @kgod

OpenStudy (anonymous):

just making sure

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

I was wondering if you could help me with a question

OpenStudy (anonymous):

of course i can, im a teacher, im here to help you guys

OpenStudy (anonymous):

ok thank you very much Mrs.Allison!!

OpenStudy (anonymous):

The graph: A coordinate plane with a colored in, closed figure. Each segment must be labeled. 2.The inequalities: Show the work for finding the inequality that defines each segment. 3.A point: Choose a point within the boundaries of the shape, and show algebraically that the point satisfies all of the inequalities.

OpenStudy (anonymous):

you are welcome, make sure to study that a little bit more and do all you homework so you can practice. @MiggityMiggityMac

OpenStudy (anonymous):

there are only those information?